# Mantis: A Simulation-Grounded Foundation Model for Disease Forecasting

### v2

@@ Discussion
Mantis demonstrates that accurate, long-range, and mechanistically interpretable infectious disease forecasting can be achieved without access to real-world training data. By learning entirely from simulated outbreaks, Mantis delivers strong performance across diverse real-world forecasting tasks and offers mechanistic explanations of its predictions through back-to-simulation attribution. These capabilities position Mantis as a practical and general-purpose tool for public health use, especially in low-resource or emerging outbreak settings where data may be limited or delayed. Because Mantis can generate forecasts with only a short initial observation window—as little as four weeks—it enables early situational awareness when conventional models are still inoperable.
While traditional forecasting models trained on historical data have been foundational for many operational systems, they often rely on disease-specific tuning and large volumes of real-world training data. Mantis departs from this paradigm: it consistently outperforms statistical, machine learning, and expert-tuned baselines across 39 real-world forecasting tasks, despite being trained on simulations alone. By proactively learning from mechanistic priors, Mantis supports plug-and-play forecasting in settings where retraining is infeasible and rapid response is essential.
The implications extend beyond public health. Mantis reflects an emerging shift toward simulation-supervised machine learning in scientific domains, where mechanistic models encode prior knowledge, and neural networks learn to generalize from them [8, 9]. This paradigm enables flexible, high-performance models that retain interpretability and domain alignment. Back-to-simulation attribution, in particular, provides a new form of interpretability tailored for science. Rather than explaining forecasts with abstract model internals, Mantis identifies which mechanistic regimes best explain the data. These attributions serve as hypotheses: falsifiable, domain-aligned, and directly actionable by public health decision-makers. This brings black-box forecasting closer to scientific reasoning—and enables the use of models not just for prediction, but for understanding.
@@
While Mantis shows broad generalization, several limitations merit consideration. As shown in [8], the model’s performance depends on the diversity and realism of training simulations, though we demonstrate generalization beyond the specific regimes directly simulated during training. Additionally, Mantis inherits any systematic biases present in the underlying mechanistic models used for training. Another limitation concerns interpretability: back-to-simulation attribution currently operates by identifying training simulations most similar to the input trajectory, and thus can only attribute to mechanisms represented in the training data. While Mantis can accurately forecast dynamics outside the training distribution, its explanations are constrained by the set of simulated mechanisms available for comparison. Expanding the breadth of simulations used specifically for attribution may enable mechanistic explanations even in out-of-distribution scenarios, since Mantis could still forecast effectively in those situations. Future work should explore active learning approaches to identify and fill critical simulation gaps, and develop methods to detect when real-world dynamics fall outside reliable prediction boundaries.
@@
Mantis illustrates a path forward for scientific forecasting systems that are interpretable, accurate, and rapidly deployable. Its out-of-the-box generalization makes it especially well-suited for early outbreak response, rural regions, and low- and middle-income countries—settings where traditional models often fail due to sparse or delayed data. By pairing simulation-based priors with flexible neural architectures, Mantis eliminates the longstanding trade-off between generalization and interpretability. It offers not just better predictions, but actionable, mechanism-level insights, making it a powerful foundation for future decision support systems in global public health.
@@ Methods
Here we provide a brief overview of the methods, with a complete description given in the Appendix.
@@
Back-to-Simulation Attribution
To generate mechanistic explanations for its forecasts, Mantis performs back-to-simulation attribution [8]. During inference, the model encodes the observed input sequence into a learned latent representation. This encoding is used to retrieve the most similar training simulations in the latent space, which are then analyzed to summarize the likely underlying mechanisms driving the forecast. Attributions are visualized by aggregating simulator parameters across top-matched simulations.
@@
CD led the study and developed the code and initial draft of the manuscript. CH and AS did the ablation studies and RM developed figure visualizations. CD, MCE, and RM all contributed substantial edits and revisions for the final manuscript.
@@ B.3 Epidemic Pattern Memory Bank
The pattern memory module (‘EpidemicPatternMemory‘) implements a learnable memory bank that stores prototypical epidemic patterns discovered during training.
@@ B.4.2 Autoregressive Generation
Nine quantiles are predicted using separate projection heads:
@@
$\quad\text{for }i\in\{1,2,\ldots,9\}$ (124)
@@ Appendix C Evaluation Data
After training, Mantis is evaluated on real-world forecasting tasks spanning multiple diseases, regions, and targets. For each task, the model is run out-of-the-box: no retraining, fine-tuning, or exposure to real data is performed. Forecast accuracy is assessed using both pointwise metrics such as mean absolute error (MAE) and probabilistic calibration metrics based on quantile coverage (Appendix D.1). Mantis’s performance is compared to established forecasting benchmarks, including the CDC Forecast Hub ensemble and individual submissions [4], as well as published models such as Chronos [15].
@@ Appendix D Additional Results
D.1 Mantis Delivers Well-Calibrated Uncertainty for Reliable Decision-Making
Accurate point forecasts alone are insufficient for public health planning—decision-makers require reliable uncertainty estimates to assess risk and allocate resources appropriately. A well-calibrated forecasting model should produce prediction intervals that contain the true values at their stated coverage rates: 50% intervals should capture observations half the time, and 90% intervals should capture observations nine times out of ten. Poor calibration undermines trust and can lead to systematic over- or under-preparation for outbreak scenarios.
We evaluated Mantis’s calibration across all diseases by comparing the predicted coverage rates of its uncertainty intervals to their empirical coverage in real-world data. For each disease and forecast horizon, we computed the proportion of observations that fell within the model’s 50% and 90% prediction intervals, then compared these empirical rates to the nominal 50% and 90% targets.
Table 5 shows that Mantis achieves excellent calibration across diverse diseases and transmission modes. Across all evaluation tasks, empirical coverage closely matches nominal coverage, with most diseases showing coverage within 2-3 percentage points of the target rates. This consistency holds across respiratory diseases (influenza, COVID-19), vector-borne diseases (dengue), and even historically eradicated pathogens (smallpox), demonstrating that simulation-based pretraining enables not only accurate point forecasts but also reliable uncertainty quantification. We visualize forecasts with both 50% and 90% CIs in figure 5.
Table 5: Prediction Interval Calibration Across Diseases
Disease 50% Coverage 90% Coverage
COVID-19 Mortality 51% 86%
COVID-19 Hospitalizations 53% 88%
Influenza Hospitalizations 51% 88%
Dengue Cases 47% 88%
Scarlet Fever 47% 88%
Hepatitis B 47% 85%
Smallpox 45% 82%
Figure 5: Forecasts with 50% CIs (dark blue shade) and 90% (light blue shade) demonstrate Mantis’s calibration.
This calibration quality is essential for operational deployment. Public health officials must trust that when Mantis indicates a 90% probability that hospitalizations will fall within a specified range, that range captures the true outcome roughly nine times out of ten. Such reliability enables evidence-based resource planning, from ventilator procurement to staffing decisions, with quantified confidence bounds that support both conservative and aggressive response strategies depending on acceptable risk levels.
The robustness of Mantis’s uncertainty estimates across diseases, time periods, and transmission modes underscores a key advantage of simulation-grounded training: by learning from diverse synthetic scenarios that span the full range of epidemiological uncertainty, the model internalizes realistic estimates of forecast confidence that transfer to real-world settings. This contrasts with models trained on limited historical data, which may exhibit over-confident predictions when encountering novel outbreak dynamics.
D.2 Ablation Studies Confirm the Importance of Scale and Mechanistic Pretraining
@@
D.2.1 Architectural Capacity Requirements
@@
D.2.2 Mechanistic Pretraining is Essential for Novel Outbreak Performance

### v3

@@ Discussion
Mantis demonstrates that accurate forecasting across diverse diseases can be achieved through training on simulated outbreaks, without access to real-world data. By learning from mechanistic simulations, Mantis delivers competitive performance across real-world forecasting tasks spanning multiple transmission modes, outbreak phases, and surveillance systems. These capabilities position Mantis as a practical tool for public health use, especially in low-resource or emerging outbreak settings where data may be limited or delayed. Because Mantis can generate forecasts with only a short initial observation window, it enables early situational awareness when conventional models require longer histories to achieve comparable accuracy. Mantis’s advantage in data-scarce settings stems from its ability to avoid overfitting to sparse initial data. Rather than extrapolating confidently from a handful of observations, Mantis produces appropriately hedged forecasts that reflect the true uncertainty across potential outbreak trajectories, enabling more reliable probabilistic planning even when historical context is minimal.
The implications extend beyond public health. Simulation-supervised learning allows mechanistic models to embed scientific knowledge into neural networks, enabling them to generalize beyond observed data [8, 9]. This paradigm offers a potential solution to a fundamental challenge in epidemiological modeling: real-world disease data are limited by privacy concerns, reporting inconsistencies, and the infrequency of major outbreaks. Simulations, by contrast, can generate unlimited training data spanning diverse scenarios. Whether simulation-based training ultimately proves superior to training on massive real-world datasets—if such data could be assembled—remains an open question. However, given current data availability constraints, the simulation-based approach offers a practical path forward for building general-purpose forecasting systems.
Mantis represents an initial attempt at a foundation model for disease forecasting, and several important questions remain unresolved. First, the optimal neural architecture for this task is unclear. While our hybrid CNN-transformer design performs well, alternative architectures may offer better performance or efficiency trade-offs. Second, the ideal balance between simulation diversity and computational cost requires further investigation. Our training set spans 400 million simulated days, generated continuously until model convergence without repeating data. Using synthetic data like this allows us to avoid the overfitting risks inherent in epoch-based training on fixed datasets. However, whether this scale is necessary or whether comparable performance could be achieved with fewer but more strategically designed simulations remains to be seen.
@@
While Mantis shows broad generalization, several limitations merit consideration. Though we demonstrate generalization beyond the specific regimes directly simulated during training, the model’s performance depends on the diversity and realism of training simulations. Additionally, Mantis inherits any systematic biases present in the underlying mechanistic models used for training. If the simulation framework systematically misrepresents certain disease dynamics—for example, categorically overestimating intervention effects—these biases will propagate to the trained model. Future work should explore approaches to identify and fill critical simulation gaps, and develop methods to validate that training simulations adequately span the space of real-world disease dynamics.
@@
Mantis illustrates a path forward for general-purpose disease forecasting systems that can be quickly deployed without disease-specific tuning. Its out-of-the-box generalization makes it especially well-suited for early outbreak response, rural regions, and low- and middle-income countries—settings where traditional models often fail due to sparse or delayed data. By pairing simulation-based priors with flexible neural architectures, this approach offers a scalable alternative to expert-tuned, disease-specific models. As a first demonstration of simulation-grounded foundation models in infectious disease forecasting, Mantis establishes both the promise and the open challenges of this paradigm for future decision support systems in global public health.
@@ Methods
A complete description of the methods is given in the Appendix.
@@
Baseline Models
We compared Mantis against two classes of baseline models: operational forecasts from the CDC COVID-19 Forecast Hub and models we re-implemented for evaluation across all diseases. For the CDC COVID-19 Forecast Hub comparison, we used forecasts as originally submitted by modeling teams during the pandemic and, to the best of our ability, evaluated both them and Mantis using the data as it was available at the time of each forecast (i.e., without any backfill or cleaning). This retrospective evaluation means our dataset contains data quality issues (e.g., missing/underreported data) that were present in the original surveillance data and likely affected all models’ performance, including Mantis. These CDC models were retrained or refitted each week to incorporate newly available data, and modeling teams could manually adjust forecasts if their automated systems produced anomalous predictions. Mantis forecasts were not modified in any way. This represents the strongest possible real-world baseline, reflecting operational forecasting practice with expert oversight under realistic data conditions.
For baselines we implemented—including statistical models (exponential smoothing and SARIMA) and deep learning models (LSTM)—we likewise retrained or refitted each model at every forecast point as new data became available. This rolling retraining procedure provides the strongest possible baseline performance for each method. Full baseline model implementation details can be found in Appendix C.
To compute relative MAE and relative WIS, we normalized all forecast errors by the performance of a naïve baseline model. This baseline generates point forecasts by projecting the most recent observed value forward to all future time points. Uncertainty intervals are constructed using the procedure from the COVIDhub-baseline model [2]: we collect all first differences from the historical time series along with their negatives (i.e., $y_{t}-y_{t-1}$ and $(y_{t}-y_{t-1})$ for all past times $t$). We form a piecewise linear approximation to the empirical cumulative distribution function of these differences, then generate 100,000 Monte Carlo samples by independently drawing sequences of $h$ differences and adding them to the most recent observed value. The resulting distribution is truncated at zero (to prevent negative forecasts) and summarized using quantiles, with the median set equal to the last observed value to eliminate sampling noise. This baseline captures the natural variability in historical changes without making assumptions about future trends, providing a neutral reference point for evaluating forecast skill.
@@
CD led the study and developed the code and initial draft of the manuscript. CH and AS did the ablation studies and RM developed figure visualizations. CD, MCE, RM, and EM all contributed substantial edits and revisions for the final manuscript.
@@ B.3 Epidemic Pattern Memory Bank
The pattern memory module (‘EpidemicPatternMemory’) implements a learnable memory bank that stores prototypical epidemic patterns discovered during training.
@@ B.4.2 Autoregressive Generation
23 quantiles are predicted using separate projection heads:
@@
$\quad\text{for }i\in\{1,2,\ldots,23\}$ (124)
@@ Appendix C Evaluation Data
After training, Mantis is evaluated on real-world forecasting tasks spanning multiple diseases, regions, and targets. For each task, the model is run out-of-the-box: no retraining, fine-tuning, or exposure to real data is performed. Forecast accuracy is assessed using both pointwise metrics such as mean absolute error (MAE), weighted interval score (WIS), and coverage metrics. Mantis’s performance is compared to established forecasting benchmarks, including the CDC Forecast Hub ensemble and individual submissions [2], as well as published models such as Chronos [18].
@@ C.3.2 Exponential Smoothing (ETS)
C.3.3 Seasonal ARIMA (SARIMA)
We implemented seasonal autoregressive integrated moving average models with automatic order selection. For each disease and forecast window, the model specification was determined using the following procedure:
  • Order selection: Auto ARIMA algorithm to determine optimal $(p,d,q)\times(P,D,Q)_{s}$ parameters based on AIC
  • Seasonality: Period determined by data frequency (e.g., 52 for weekly)
  • Training procedure: Model refit at each time step using all available historical data up to the forecast origin
  • Forecast generation: Rolling forecasts with the refitted model producing multi-step-ahead predictions
  • Model selection: Best SARIMA specification selected per disease based on AIC performance during the initial training period
C.3.4 Long Short-Term Memory Network (LSTM)
We implemented a deep recurrent neural network baseline using a multi-layer LSTM architecture. The model was retrained at each forecast point to incorporate all available historical data:
  • Architecture: 6-layer LSTM with hidden dimension of 128 units per layer
  • Regularization: Dropout rate of 0.5 applied between LSTM layers
  • Data preprocessing: Standard scaler applied to normalize input time series
  • Training procedure: Model retrained from scratch at each time step using all available historical data
  • Training duration: 50 epochs per retraining with early stopping based on validation loss
  • Forecast generation: Rolling forecasts produced after retraining at each forecast origin
@@ Appendix D Additional Results
D.1 Ablation Studies Confirm the Importance of Scale and Mechanistic Pretraining
@@
D.1.1 Architectural Capacity Requirements
@@
D.1.2 Mechanistic Pretraining is Essential for Novel Outbreak Performance
@@
D.2 Performance Consistency Across Population Scales
To assess whether Mantis’s performance varies systematically with jurisdiction size, we analyzed the relationship between state population and forecast accuracy for the COVID-19 mortality forecasting task. We computed the mean relative MAE for each U.S. state and territory across all forecast dates in the evaluation period (April 2020 through November 2021) and examined the correlation with 2020 Census population estimates.
Figure 4 shows Mantis’s relative MAE versus state population on a logarithmic scale. We excluded Vermont from this analysis due to anomalous worse performance, perhaps due to data quality issues: in Vermont’s COVID-19 surveillance data, hospitalizations significantly lag deaths (rather than preceding them as expected), suggesting potential reporting artifacts that confound model evaluation. Among the remaining 51 jurisdictions, we observed a weak negative correlation between log population and relative MAE ($R^{2}=0.06$), meaning Mantis performs slightly better in larger states. However, the low $R^{2}$ value suggests that Mantis performance is relatively consistent across diverse population scales.
This consistency is positive given that smaller jurisdictions present distinct forecasting challenges due to noisier case trajectories and sometimes less available data. The weak relationship between population and performance suggests that Mantis’s simulation-based training enables robust forecasting across the spectrum of population sizes from states like Wyoming to California.
Figure 4: Mantis delivers consistent performance across population scales. Relative MAE versus state population for COVID-19 mortality forecasts across 51 U.S. states and territories (Vermont excluded as an outlier). Each point represents the mean relative MAE for a jurisdiction across all forecast dates from April 2020 through November 2021. Population is shown on a logarithmic scale (2020 Census). A weak negative correlation ($R^{2}=0.06$) indicates slightly better performance in larger states, but the low coefficient of determination suggests that Mantis achieves relatively uniform accuracy across diverse population sizes.